# Rapid Dissipative Ground State Preparation at Chemical Transition States

### v1

@@ C.2 Per-segment cost of dissipative continuation
Our algorithm uses a short-time dissipative primitive (a filtered-jump Lindbladian step) at geometry $s_{i+1}$ initialized by a warm start transported from $s_{i}$ (e.g., via Procrustes-aligned orbital rotations plus state reuse). Let $\varepsilon$ denote the target failure probability for cooling at that segment. Under the conditions of Lemmas 5 and 6 provide a per-segment mixing time
@@ Appendix D Phase Randomization and Dynamic Cooling versus Dissipative Evolution
Phase randomization (PR) [31] and dynamic cooling (DC) [34] provide unitary-centric alternatives to using an explicit Lindbladian cooling primitive. Both are useful comparators in the reaction-path setting because they exploit the same structural feature as dissipative continuation: a warm start at $s_{i}$ is already concentrated in the low-energy subspace of $H(s_{i+1})$ when the path is discretized finely enough. The purpose of this section is to justify the gate-complexity scalings used in Table 1, highlighting how the dependence on $\|H\|$, $\Delta_{\min}$, and $N_{o}$ arises from (i) Hamiltonian simulation costs, (ii) the time scale required to resolve gaps, and (iii) the number of randomization / cooling iterations required at each path point.
@@ D.2 Phase randomization
The PR traversal algorithm [31] stabilizes transport along the discretized path by using this dephasing as an approximate projective step. The relevant cost contributions are: (i) the cost per randomized evolution, (ii) the characteristic time scale required to resolve the gap, and (iii) the number of PR attempts (or equivalently, the repetition overhead) needed to achieve an overall success probability $p_{\mathrm{succ}}$.
@@ D.3 Dynamic cooling
Dynamic cooling alternates a dephasing/phase-mixing step with an explicitly energy-selective unitary procedure, implemented using QSP/QSVT transformations interleaved with a weak perturbation that induces transitions biased toward lower energies [34]. The scaling used in Table 1 can be recovered by combining (i) the same Hamiltonian-simulation cost model of Eq. 92, (ii) the same DK-based discretization $N_{H}=\Theta(C_{\mathrm{DK}}^{2}/\epsilon_{E})$, and (iii) the DC iteration complexity at each geometry. In the DC analysis, the number of calls to the underlying $e^{-iHt}$ primitive at a fixed geometry is governed by an effective dimension parameter $d$ that controls the schedule/energy-resolution regime of the procedure. In the regime relevant to reaction-path electronic-structure instances, one takes
@@ Appendix E Ground-state overlap from occupied-space principal angles
    with $\alpha$ depending only on the Gevrey class [20].
@@
The Gevrey part is an existence statement: [20] constructs Gevrey-class filters that approximate the one-sided step outside a spectral margin while keeping the time-domain kernel essentially supported on a window of length $O\!\big(\frac{\|H\|}{\Delta}[\log(1/\varepsilon_{\mathrm{leak}})]^{\alpha}\big)$. Applying the same reduction Eq. 113 yields the leakage guarantee. ∎
@@ Appendix G Reducing the Cost of Simulating Dissipative Dynamics via qDRIFT-inspired Jump Sampling
This randomized splitting is the open-system analouge of the qDRIFT protocol for Hamiltonian simulation [62]. Note that by a Taylor expansion,
@@
Following [62], the resulting error can be characterized in the diamond norm and scales as
@@ Appendix H Logical Quantum Circuit Construction for Dissipative Ground State Preparation
where $\{K_{a}\}$ are the filtered jump operators defined in Definitions 5 and 33 and we implement a circuit that realizes a CPTP map $e^{\tau\mathcal{L}_{a}}$.
@@
In practice, we can reduce the required quantum resources to implement the dissipative evolution over the entire jump set $\mathcal{A}$ by randomly sampling jump operators for each time step $e^{\tau\mathcal{L}_{a}}$ where $A_{a}$ are drawn i.i.d.. Ding et. al. [20] provide an efficient circuit for implementing the dissipative evolution of a single filtered jump operator $K_{a}$ that only requires a single ancilla in addition to the system register (excluding ancilla for block encodings). The operator $W(\sqrt{\tau})$ acting on $n$ system qubits and one ancilla qubit that approximates a time step of dissipative evolution for the target Lindbladian $\mathcal{L}$ to second order,
@@
Our circuit construction is inspired by the Monte Carlo–style quantum algorithm for ground-state preparation in the early fault-tolerant regime [20], which features a Lindbladian whose target state is stationary. In the paper, the authors design a single-ancilla circuit that simulates a Lindblad dynamics step by splitting the evolution into a purely Hamiltonian part and a purely dissipative part with first-order trotter method with just one extra qubit. Each block is a short interaction between the ancilla and a simple system operator, with the ancilla axis encoding the complex filter coefficient, so the whole channel is built from a sequence of elementary ancilla rotations and controlled interactions. We adapt and extend this in our own circuit design along the reaction path.
@@
Therefore, we seek an efficient quantum circuit implementation of the dissipative update $e^{\tau\mathcal{L}_{a}}$ at each time step. We can follow the recent work [63] in which the authors show that Stinespring dilation can be used to implement dissipative evolution via Hamiltonian simulation and measurement.
@@
To implement the jump operator which governs the dissipation part of the Lindbladian Master equation, we follow the structure of [20] by introducing single ancilla qubit for the environment as below
@@
From Eqs. 136 and 141, we can implement the circuit as following
@@ Appendix I Logical Quantum Resources Estimation for Dissipative Ground State Preparation
Table 2: Qubit counts and Toffoli costs per single dissipative time step $e^{\tau\mathcal{L}(s)}$, implemented by a second-order Strang splitting operator $W(\sqrt{\tau})$ with $N_{\omega}=1000$ Fourier modes for the filter function, together with QPE Toffoli counts (for $\sigma_{\mathrm{PEA}}=1$ mHa) using DFTHC+BLISS+SA block encodings for systems [44].

### v2

@@ C.2 Per-segment cost of dissipative continuation
Our algorithm uses a short-time dissipative primitive (a filtered-jump Lindbladian step) at geometry $s_{i+1}$ initialized by a warm start transported from $s_{i}$ (e.g., via Procrustes-aligned orbital rotations plus state reuse). Let $\varepsilon$ denote the target failure probability for cooling at that segment. Under the conditions of Lemmas 6 and 7 provide a per-segment mixing time
@@ Appendix D Phase Randomization and Dynamic Cooling versus Dissipative Evolution
Phase randomization (PR) [Boixo2009Eigenpath] and dynamic cooling (DC) [motlagh2024_cooling] provide unitary-centric alternatives to using an explicit Lindbladian cooling primitive. Both are useful comparators in the reaction-path setting because they exploit the same structural feature as dissipative continuation: a warm start at $s_{i}$ is already concentrated in the low-energy subspace of $H(s_{i+1})$ when the path is discretized finely enough. The purpose of this section is to justify the gate-complexity scalings used in Table 1, highlighting how the dependence on $\|H\|$, $\Delta_{\min}$, and $N_{o}$ arises from (i) Hamiltonian simulation costs, (ii) the time scale required to resolve gaps, and (iii) the number of randomization / cooling iterations required at each path point.
@@ D.2 Phase randomization
The PR traversal algorithm [Boixo2009Eigenpath] stabilizes transport along the discretized path by using this dephasing as an approximate projective step. The relevant cost contributions are: (i) the cost per randomized evolution, (ii) the characteristic time scale required to resolve the gap, and (iii) the number of PR attempts (or equivalently, the repetition overhead) needed to achieve an overall success probability $p_{\mathrm{succ}}$.
@@ D.3 Dynamic cooling
Dynamic cooling alternates a dephasing/phase-mixing step with an explicitly energy-selective unitary procedure, implemented using QSP/QSVT transformations interleaved with a weak perturbation that induces transitions biased toward lower energies [motlagh2024_cooling]. The scaling used in Table 1 can be recovered by combining (i) the same Hamiltonian-simulation cost model of Eq. 94, (ii) the same DK-based discretization $N_{H}=\Theta(C_{\mathrm{DK}}^{2}/\epsilon_{E})$, and (iii) the DC iteration complexity at each geometry. In the DC analysis, the number of calls to the underlying $e^{-iHt}$ primitive at a fixed geometry is governed by an effective dimension parameter $d$ that controls the schedule/energy-resolution regime of the procedure. In the regime relevant to reaction-path electronic-structure instances, one takes
@@ Appendix E Ground-state overlap from occupied-space principal angles
    with $\alpha$ depending only on the Gevrey class [ding2024single].
@@
The Gevrey part is an existence statement: [ding2024single] constructs Gevrey-class filters that approximate the one-sided step outside a spectral margin while keeping the time-domain kernel essentially supported on a window of length $O\!\big(\frac{\|H\|}{\Delta}[\log(1/\varepsilon_{\mathrm{leak}})]^{\alpha}\big)$. Applying the same reduction Eq. 115 yields the leakage guarantee. ∎
@@ Appendix G Reducing the Cost of Simulating Dissipative Dynamics via qDRIFT-inspired Jump Sampling
This randomized splitting is the open-system analouge of the qDRIFT protocol for Hamiltonian simulation [Campbell2019]. Note that by a Taylor expansion,
@@
Following [Campbell2019], the resulting error can be characterized in the diamond norm and scales as
@@ Appendix H Logical Quantum Circuit Construction for Dissipative Ground State Preparation
where $\{K_{a}\}$ are the filtered jump operators defined in Definitions 6 and 35 and we implement a circuit that realizes a CPTP map $e^{\tau\mathcal{L}_{a}}$.
@@
In practice, we can reduce the required quantum resources to implement the dissipative evolution over the entire jump set $\mathcal{A}$ by randomly sampling jump operators for each time step $e^{\tau\mathcal{L}_{a}}$ where $A_{a}$ are drawn i.i.d.. Ding et. al. [ding2024single] provide an efficient circuit for implementing the dissipative evolution of a single filtered jump operator $K_{a}$ that only requires a single ancilla in addition to the system register (excluding ancilla for block encodings). The operator $W(\sqrt{\tau})$ acting on $n$ system qubits and one ancilla qubit that approximates a time step of dissipative evolution for the target Lindbladian $\mathcal{L}$ to second order,
@@
Our circuit construction is inspired by the Monte Carlo–style quantum algorithm for ground-state preparation in the early fault-tolerant regime [ding2024single], which features a Lindbladian whose target state is stationary. In the paper, the authors design a single-ancilla circuit that simulates a Lindblad dynamics step by splitting the evolution into a purely Hamiltonian part and a purely dissipative part with first-order trotter method with just one extra qubit. Each block is a short interaction between the ancilla and a simple system operator, with the ancilla axis encoding the complex filter coefficient, so the whole channel is built from a sequence of elementary ancilla rotations and controlled interactions. We adapt and extend this in our own circuit design along the reaction path.
@@
Therefore, we seek an efficient quantum circuit implementation of the dissipative update $e^{\tau\mathcal{L}_{a}}$ at each time step. We can follow the recent work [Ding2024simulating] in which the authors show that Stinespring dilation can be used to implement dissipative evolution via Hamiltonian simulation and measurement.
@@
To implement the jump operator which governs the dissipation part of the Lindbladian Master equation, we follow the structure of [ding2024single] by introducing single ancilla qubit for the environment as below
@@
From Eqs. 138 and 143, we can implement the circuit as following
@@ Appendix I Logical Quantum Resources Estimation for Dissipative Ground State Preparation
Table 2: Qubit counts and Toffoli costs per single dissipative time step $e^{\tau\mathcal{L}(s)}$, implemented by a second-order Strang splitting operator $W(\sqrt{\tau})$ with $N_{\omega}=1000$ Fourier modes for the filter function, together with QPE Toffoli counts (for $\sigma_{\mathrm{PEA}}=1$ mHa) using DFTHC+BLISS+SA block encodings for systems [Low_2025].
Appendix J Reaction path optimization
This appendix records a simple but useful fact: the Davis–Kahan integral
$C_{\mathrm{DK}}=\int_{0}^{1}\frac{\|\partial_{s}H(s)\|_{2}}{\Delta(s)}\,ds$
is controlled by a weighted path-length functional of the nuclear trajectory. This provides a direct route for connecting a reaction-path optimization objective (such as the one implemented in our N3 example) to reductions in $C_{\mathrm{DK}}$, and therefore to reductions in the discretization cost $N_{H}$ in Theorem 3.
Lemma 8 (Chain-rule bound for the DK integrand).
Let $\mathbf{R}:[0,1]\to\mathbb{R}^{3N}$ be a $C^{1}$ reaction path and define $H(s)=H_{\mathrm{el}}(\mathbf{R}(s))$. Assume $H_{\mathrm{el}}$ is Fréchet differentiable in $\mathbf{R}$ along the image of the path, with derivative $D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R}):\mathbb{R}^{3N}\to\mathsf{Herm}$. Define the operator-Lipschitz factor
$L_{H}(\mathbf{R})\coloneqq\sup_{\|v\|=1}\bigl\|D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R})[v]\bigr\|_{2}.$
Then for all $s\in[0,1]$,
$\frac{\|\partial_{s}H(s)\|_{2}}{\Delta(s)}\leq\frac{L_{H}(\mathbf{R}(s))}{\Delta(s)}\,\|\mathbf{R}^{\prime}(s)\|.$ (152)
Consequently,
$C_{\mathrm{DK}}\leq\int_{0}^{1}\frac{L_{H}(\mathbf{R}(s))}{\Delta(s)}\,\|\mathbf{R}^{\prime}(s)\|\,ds.$ (153)
Proof.
Let $s\in[0,1]$ be fixed. By the chain rule in Banach spaces,
$\partial_{s}H(s)=D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R}(s))[\mathbf{R}^{\prime}(s)].$
Using the definition of $L_{H}(\mathbf{R}(s))$ and homogeneity in the direction argument,
$\|\partial_{s}H(s)\|_{2}=\bigl\|D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R}(s))[\mathbf{R}^{\prime}(s)]\bigr\|_{2}\leq L_{H}(\mathbf{R}(s))\,\|\mathbf{R}^{\prime}(s)\|.$
Dividing by $\Delta(s)>0$ gives Eq. 152. Integrating both sides from 0 to $1$ yields Eq. 153. ∎
Corollary 1 (Uniform bound and a clean geometry surrogate).
Suppose that along the path image one has
$L_{H}(\mathbf{R}(s))\leq L_{\max}\qquad\text{and}\qquad\Delta(s)\geq\Delta_{\min}>0\qquad\text{for all }s\in[0,1].$
Then
$C_{\mathrm{DK}}\leq\frac{L_{\max}}{\Delta_{\min}}\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|\,ds.$ (154)
In particular, if $\mathbf{R}$ is parameterized by arc-length (so $\|\mathbf{R}^{\prime}(s)\|$ is constant), reducing the path length $\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|ds$ reduces $C_{\mathrm{DK}}$ by the same factor.
Proof.
Apply Eq. 153 and upper bound the numerator and lower bound the denominator uniformly:
$C_{\mathrm{DK}}\leq\int_{0}^{1}\frac{L_{H}(\mathbf{R}(s))}{\Delta(s)}\|\mathbf{R}^{\prime}(s)\|ds\leq\frac{L_{\max}}{\Delta_{\min}}\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|ds.$
Appendix K Dissipative Evolution Algorithm Pseudocode
1
Input: Endpoint geometries $\mathbf{R}_{\mathrm{R}},\mathbf{R}_{\mathrm{P}}$, target geometry $s^{\star}$, target energy error $\epsilon_{E}$, jump set $\mathcal{A}$, filter function $\hat{f}(\omega)$
Output: Approximate transition-state ground state $\rho^{\star}$ and energy estimate $E^{\star}$
2
3Choose an initial reaction path $\mathbf{R}_{0}(s)$ connecting $\mathbf{R}_{\mathrm{R}}$ and $\mathbf{R}_{\mathrm{P}}$;
4
5Optimize the path by minimizing
$J(\mathbf{R})=\lambda_{\mathrm{len}}\!\int_{0}^{1}\|\mathbf{R}^{\prime}(s)\|\,ds+\lambda_{\mathrm{curv}}\!\int_{0}^{1}\|\mathbf{R}^{\prime\prime}(s)\|^{2}\,ds+\lambda_{\mathrm{CI}}\!\int_{0}^{1}\Phi_{\mathrm{CI}}(\mathbf{R}(s))\,ds$
to obtain a smooth path $\mathbf{R}(s)$ that avoids sharp projector rotation and small-gap regions;
6
7Discretize the optimized path into $\{s_{i}\}_{i=0}^{N_{H}-1}$ with $s_{0}=0$ and $s_{N_{H}-1}=s^{\star}$;
8
9For each $s_{i}$, build the electronic Hamiltonian
$H_{i}=H_{\mathrm{el}}(\mathbf{R}(s_{i}))$
and align neighboring orbital spaces to maintain a smooth gauge along the path;
10
11Prepare an approximate ground state $\rho_{0}$ for $H_{0}$ using a tractable classical or hybrid method (e.g., DMRG + dissipative cooling) ;
12
13for $i=0$ to $N_{H}-2$ do
14   Construct filtered jump operators $\{K_{a}^{(i+1)}\}_{a\in\mathcal{A}}$ for $H_{i+1}$;
15
16  Starting from $\rho_{i}$, apply the dissipative primitive for $t_{i}$ steps to obtain
$\rho_{i+1}\approx\left(e^{\tau\mathcal{L}_{i+1}}\right)^{t_{i}}(\rho_{i}),$
where
$\mathcal{L}_{i+1}[\cdot]=\sum_{a\in\mathcal{A}}K_{a}^{(i+1)}(\cdot)K_{a}^{(i+1)\dagger}-\frac{1}{2}\left\{K_{a}^{(i+1)\dagger}K_{a}^{(i+1)},\cdot\right\}.$
17 end for
18
19Set $\rho^{\star}\leftarrow\rho_{N_{H}-1}$ and estimate
[…]
20return $(\rho^{\star},E^{\star})$;
21
Algorithm 1 Dissipative evolution with reaction-path optimization